1. If you lend $3600 to a friend for 15 months at 10% annual simple interest, find the future value of the loan.
$_________
2. What are the future value and the interest earned if $3400 is invested for 7 years at 8% compounded quarterly? (Round your answers to the nearest cent.)
future value $_________
interest earned $_________
3. How long (in years) would $400 have to be invested at 7%, compounded continuously, to amount to $905? (Round your answer to the nearest whole number.)
years__________

Answer by stanbon(75887)

1. If you lend $3600 to a friend for 15 months at 10% annual simple interest, find the future value of the loan.
FV = PV x (1+r)^n
FV = 3600 x (1.1)^(15/12)
FV = $4055.49_________
2. What are the future value and the interest earned if $3400 is invested for 7 years at 8% compounded quarterly? (Round your answers to the nearest cent.)
future value $3400 x (1.08)^(7*4) = $29332.16_________
interest earned $29332.16-3400 = $25932.16
3. How long (in years) would $400 have to be invested at 7%, compounded continuously, to amount to $905? (Round your answer to the nearest whole number.)
A(t) = P*e^(rt)
905 = 400*e^(0.07t)
2.26 = e^(0.075)
0.075t = ln(2.26)
0.075t = 0.8154
t = 11 when rounded up
years = 11 when rounded up
_________
Cheers,
Stan H.
----------